# On computing and the complexity of computing higher-order $U$-statistics, exactly

### v1

@@ B.1 Preparatory Materials
Lemma 8 (Möbius Inversion Formula, Dual Form, (stanley2011enumerative)).
@@ B.2 Proof of Lemma 5
Appendix C Proof of Observation 1 and Results in Section 4.1
C.1 Proof of Observation 1
@@
Lemma 9 (Bounds on Treewidth).
@@
Proof.
The first inequality $D(G)\leq\mathsf{tw}(G)$ is established in Lemma 11 of bodlaender2011treewidth.
The second inequality is stated as Lemma 8 in the same reference.
The final claim follows directly from Definition 7: treewidth is defined as the minimum over all possible vertex elimination orderings, whereas the right-hand side corresponds to the class of orderings in which the vertex $v$ is eliminated first. ∎
Lemma 10 (Maximum Treewidth under Vertex and Edge Constraints).
@@
    Support set: For any fixed $e\in\mathbb{N}^{+}$, $\mathcal{G}_{n,e}$ is nonempty iff $n\in\mathrm{n}(e)\coloneqq\left\{n\in\mathbb{N}^{+}\,\middle|\,\binom{n}{2}\geq e,n\leq 2e\right\}$.
@@
    Complete-graph treewidth: $\mathrm{tt}(n,\binom{n}{2})=n-1$, $\mathrm{tt}(n,\binom{n}{2}-1)=n-2$.
@@
      If $1\leq\frac{2e}{n}<2$, then $\mathrm{tt}(n,e)\leq\max\{1,\mathrm{t}(e-1)\}$;
@@
      If $1\leq\frac{2e}{n}<3$, then $\mathrm{tt}(n,e)\leq\max\{2,\mathrm{t}(e-1)\}$;
@@
C.2 Proof of Results in the Rightmost Column of Table 1
We can derive the sharp treewidth bounds presented in Table 1 using similar techniques to the above. For each $m\in\{2,3,\cdots,12\}$, the corresponding decomposition signature takes the following form: ${\mathcal{A}}_{m}=((1,2),\cdots,(m-1,m))$. Hence by Definition 10, the corresponding decomposition graph $G_{{\mathcal{A}}_{m}}$ is the graph with $V(G_{{\mathcal{A}}_{m}})=[m]$, $E(G_{{\mathcal{A}}_{m}})=\{\{i,i+1\}\}_{i=1}^{m-1}$, whose structure is illustrated in Figure 4(a).
@@ C.2 Proof of Results in the Rightmost Column of Table 1
Table 7: Upper bounds of $M(m)$ at different $m$’s.
@@
Table 8: Lowers bound of $M(m)$: the exemplified graph whose treewidth reaches the lower bound of $M(m)$ at different $m$’s.
@@
To this end, we will first establish a lemma (Lemma 11) stating that a simple graph with at most $6$ edges has treewidth $3$ if and only if it is $K_{4}$ (i.e. the complete graph with $4$ vertices). Then, another lemma (Lemma 12) will show that the set of all quotient graphs of $G_{{\mathcal{A}}_{7}}$ does not contain $K_{4}$. The proof is now complete. ∎
@@
This map can be interpreted as a coloring of the vertices: each vertex $k$ is colored by the index of the subset $B_{i}$ which it belongs to, so that all vertices in $Q_{i}$ are dyed by color $i$. By above analysis, each pair of adjacent vertices is assigned different colors, thus for each $i\in[6]$, $\{Q_{c(i)},Q_{c(i+1)}\}$ is an edge in $G_{{\mathcal{A}}_{7}}/\pi^{*}$.
Then, the sequence of edges
$\{Q_{c(1)},Q_{c(2)}\},\ \{Q_{c(2)},Q_{c(3)}\},\ \{Q_{c(3)},Q_{c(4)}\},\ \{Q_{c(4)},Q_{c(5)}\},\ \{Q_{c(5)},Q_{c(6)}\},\ \{Q_{c(6)},Q_{c(7)}\}$
corresponds to an Euler trail (i.e., a sequence of edges in which consecutive edges share a common endpoint, visiting every edge of the graph exactly once without repetition.) in the quotient graph $G_{{\mathcal{A}}_{7}}/\pi^{*}=K_{4}$.

### v2

@@ 5 Concluding Remarks
Lemma 8 (Möbius Inversion Formula, Dual Form, stanley2011enumerative).
@@
[…]
In this section, we present an open problem that we ourselves find intriguing. As we indicated in Remark 6 at the end of Section 3, the treewidth of the decomposition graph associated with the Einsum notation of the underlying $V$-statistic relates to its time complexity. But the treewidth of a graph is not easy to compute either analytically (harvey2018treewidth) or numerically (arnborg1987complexity). It will be interesting to get a sharp bound on treewidth using the number of vertices or edges of the graph. As mentioned in Remark 6, kneis2009bound has shown that the treewidth is bounded at most linearly by the edge size, or more precisely bounded by $e/5.769+O(\log n)$ where $n\coloneqq|V|$ denotes the number of vertices in $G$. However, we do not know whether this bound can be further improved. Below we provide the maximum tree-width that is possible for a graph with edge size no greater than $15$, to hopefully serve as a starting point of fully addressing this problem.
Observation 1 (Treewidth and Edge Count).
Let
[…]
@@
We first present some results on inequalities involving treewidth. The first two inequalities are from bodlaender2011treewidth, while the last one follows directly from Definition 7.
Lemma 9.
@@
We then present results concerning the treewidth of graphs subject to constraints on the number of edges or vertices.
Lemma 10.
@@
    Support set: For any fixed $e\in\mathbb{N}^{+}$, $\mathcal{G}_{n,e}$ is nonempty if and only if
    $n\in\mathrm{n}(e)\coloneqq\left\{n\in\mathbb{N}^{+}\,\middle|\,\binom{n}{2}\geq e,n\leq 2e\right\}.$
@@
    Treewidth of complete graph: $\mathrm{tt}(n,\binom{n}{2})=n-1$, $\mathrm{tt}(n,\binom{n}{2}-1)=n-2$.
@@
      If $1\leq\frac{2e}{n}<2$, then $\mathrm{tt}(n,e)\leq\max\{1,\mathrm{t}(e-1)\}$.
@@
      If $1\leq\frac{2e}{n}<3$, then $\mathrm{tt}(n,e)\leq\max\{2,\mathrm{t}(e-1)\}$.
@@
[…]
The following table lists the total number of $V$-statistics and the number of floating-point operations (FLOPs) that need to be computed before and after applying the sparsification trick when the sample size is $10^{4}$. The rightmost column of Table 4 also records the value of $M$ defined in Corollary 2.
Table 4: Exact number of floating-point operations (FLOPs), $V$-statistic decompositions, intermediate tensor sizes, and the maximum treewidth $M$ across all decomposition graphs for HOIF orders from $2$ to $12$, with and without the sparsification trick, where the Einsum ordering is found by applying opt-einsum with optimization scheme greedy. Sample size is fixed at $n=10000$.
Order ($m$) V-stat terms (Bell number) V-stat terms (sparsification) FLOPs (initial) FLOPs (sparsification) $M$
2 2 1 $2.00020\times 10^{8}$ $2.00000\times 10^{8}$ 11
3 5 2 $060\times 10^{12} 2.00020$$020\times 10^{12}$ 11
4 15 5 $8.00310\times 10^{12}$ $4.00150\times 10^{12}$ 22
5 52 15 $5.20133\times 10^{13}$ $60045\times 10^{13} 22$
6 203 52 $2.94069\times 10^{14}$ $1.20020\times 10^{14}$ 22
7 877 203 $1.78236\times 10^{15}$ $6.32091\times 10^{14}$ 22
8 4140 877 $31054\times 10^{17} 2.2$$23430\times 10^{17}$ 33
9 21147 4140 $5.29126\times 10^{18}$ $47959\times 10^{18} 33$
10 115975 21147 $2.07990\times 10^{21}$ $2.04048\times 10^{21}$ 33
11 678570 115975 $5.50026\times 10^{22}$ $3.72144\times 10^{22}$ 44
12 4213597 678570 $1.13401\times 10^{24}$ $6.62181\times 10^{23}$ 44
[…]
@@
Table 5: Upper bounds of $M(m)$ for HOIFs: via $M(m)\leq\mathrm{t}(m-1)$.
@@
Table 6: Lowers bound of $M(m)$ for HOIFs: via finding the exemplified graph.
@@
To this end, we will first establish a lemma (Lemma 11) stating that a simple graph with at most $6$ edges has treewidth $3$ if and only if it is $K_{4}$ (i.e. the complete graph with $4$ vertices). Then, another lemma (Lemma 12) will show that the set of all quotient graphs of $G_{\mathcal{A}_{7}}$ does not contain $K_{4}$. The proof is then complete. ∎
@@
By above analysis, all 6 edges in $G_{\mathcal{A}_{7}}/\pi^{*}$ must be
$\{Q_{c(1)},Q_{c(2)}\},\ \{Q_{c(2)},Q_{c(3)}\},\ \{Q_{c(3)},Q_{c(4)}\},\ \{Q_{c(4)},Q_{c(5)}\},\ \{Q_{c(5)},Q_{c(6)}\},\ \{Q_{c(6)},Q_{c(7)}\},$
which forms an Euler trail (i.e., a sequence of edges in which consecutive edges share a common endpoint, visiting every edge of the graph exactly once without repetition.) in the quotient graph $G_{\mathcal{A}_{7}}/\pi^{*}=K_{4}$.
@@
[…]
In this section, we first present the $U$-statistic representations for all 3-vertex and 4-vertex motifs mentioned in Section 4.2. Letting $B\equiv 1-A$, the corresponding $U$-statistics for motif counts for specific $3$-vertex motifs and $4$-vertex motifs (see Figure 5 for which motifs $\mathsf{r}_{1}$ to $\mathsf{r}_{8}$ encode) read as follows:
(a) $\mathsf{r}_{1}$: V-shape
(b) $\mathsf{r}_{2}$: Triangle
(c) $\mathsf{r}_{3}$ : 3-star
(d) $\mathsf{r}_{4}$ : Fork
(e) $\mathsf{r}_{5}$ : Tailed triangle
(f) $\mathsf{r}_{6}$: Square
(g) $\mathsf{r}_{7}$: House
(h) $\mathsf{r}_{8}$ : 4-clique
Therefore, motif counts for $m$-vertex motifs naturally correspond to a complete decomposition signature $\mathcal{A}_{\mathsf{Motif},m}=((s,t))_{1\leq s<t\leq m}$, where the associated decomposition graph $G_{\mathcal{A}_{\mathsf{Motif},m}}$ is the complete graph $K_{m}$. This structure has two important implications.
First, by leveraging our sparsification technique (Lemma 2 and Remark 4), the computation reduces to a single dominant term, i.e.,
$\Pi_{m}^{\mathcal{A}_{\mathsf{Motif},m}}=\{\pi_{m}\},\quad\text{where }\pi_{m}\coloneqq\big\{\{1\},\{2\},\ldots,\{m\}\big\}.$
Second, this also reveals a fundamental limitation. Even after sparsification, the resulting term still induces the same complete decomposition structure,
$G_{\mathcal{A}_{\mathsf{Motif},m}}/\pi_{m}=K_{m},$
and therefore the computational complexity of exact motif counting cannot be reduced below $O(n^{m})$ within our framework.
We next present the experimental results for 4-vertex motif counting and GPU-based triangle counting, which are summarized in Table 7 and Table 8, respectively. The overall message is quite similar to that in Table 2 presented in Section 4.2.
Table 7: Runtime comparison of exact all $4$-vertex motif counts using u-stats, Peregrine and igraph on Erdős–Rényi graphs $G(n,p)$ with $n=2000$. Experiments were run on Intel Xeon Scalable Cascade Lake 6248 CPUs (2.5GHz, 40 total cores) with memory of 192 GB. “OOT” denotes instances that exceeded the time limit of 3600 seconds.
Edge Prob. $p$ u-stats Time (s) Peregrine Time (s) igraph Time (s)
0.0005 3.8413 0.1024 0.0752
0.0010 18.7463 0.1141 0.0873
0.0050 41.0730 0.1132 0.2645
0.0100 40.7701 0.1116 1.7099
0.0500 41.5188 0.2079 610.0180
0.1000 40.5648 0.8243 OOT
0.2000 41.2671 5.8045 OOT
0.4000 43.3735 48.7363 OOT
0.6000 39.6022 159.4699 OOT
0.8000 40.6578 292.4395 OOT
Table 8: Runtime comparison of exact triangle counting using u-stats and cuGraph on Erdős–Rényi graphs $G(n,p)$ with $n=10000$. Experiments were run on a single GPU (NVIDIA RTX 4090, 24GB) “OOM” indicates out-of-memory.
Edge Prob. $p$ u-stats Time (s) cuGraph Time (s)
0.001 2.8014 0.1536
0.005 2.9564 0.5664
0.010 3.3363 1.0379
0.020 3.7830 1.9797
0.050 5.0856 4.9804
0.080 6.3764 8.1346
0.100 7.2829 10.3661
0.150 9.6873 15.3828
0.200 11.7283 OOM
0.800 37.6880 OOM